OpenStudy (anonymous):

55% of what number is 70

OpenStudy (anonymous):

\[number=\frac{ 70 }{ 0.55 } =127.27(27)\]

OpenStudy (dangerousjesse):

is/of = %/100 70/x = 55/100 55x = 7000 x = 7000/55

OpenStudy (jgirl128):

is/of = %/100 70/x = 55/100 55x = 7,000 x = 7000/55

OpenStudy (dangerousjesse):

openstudy.com/updates/4f4804f2e4b00c3c5d3271e3

OpenStudy (anonymous):

127.27(27)=1400/11

OpenStudy (anonymous):

x=7000/55 because 70/x=55/100

OpenStudy (jgirl128):

@Syderitic rather than giving them the answer right away, let them try to figure it out. :) just so they understand

OpenStudy (anonymous):

@jgirl128 i'm doing this or mahjong...right now, it's this. I do provide full answer...i think.

OpenStudy (dangerousjesse):

@jgirl128 Rather than stealing another user's answer, try giving whomever actually write that answer credit.

OpenStudy (dangerousjesse):

wrote*

OpenStudy (anonymous):

I did the math...

OpenStudy (dangerousjesse):

Thank you for working it out @INeedHelpWithThisQuestion